OpenStudy (anonymous):

Solve the equation (4x + 1)(3x - 2) = 0

OpenStudy (anonymous):

if the product is 0 then either 4x+1 = 0 or 3x-2 = 0 so, 4x+1=0 4x=-1 x= -1/4 and 3x-2=0 3x=2 x=2/3

OpenStudy (anonymous):

either 4x+1 or 3x-2 have to be 0, since 0 multiplied with something = 0. So 4x=-1 and 3x-2 = 2. Then x= -1/4 or 3/2

OpenStudy (earthcitizen):

\[x = -1/4, x = 2/3\]\[12 x^2-5 x-2 = 0\]